In this talk we discuss a construction of indecomposable p-algebras with exponent p and large index. This is closely connected with the study of degeneracy in matrices defining abelian crossed products and torsion in CH2 of Severi-Brauer varieties as developed by Karpenko and others. These connections are not completely understood and some open questions will be discussed.